# Suppression of Bogoliubov momentum pairing and emergence of non-Gaussian correlations in ultracold interacting Bose gases

### v1

@@ II Extracting two-body connected correlations
This generic equation is valid for both Gaussian and non-Gaussian states.
@@
IV Correlation functions in momentum space and their scaling behavior
IV.1 Definitions of the correlation functions
In this work, we consider the following connected correlations of the (momentum) density operator $N(\bm{k})$ (defined as $N(\bm{k})=a^{\dagger}_{\bm{k}}a_{\bm{k}}$),
$G_{c}^{(2)}(\bm{k}_{1},\bm{k}_{2})$ $=\langle N(\bm{k}_{1})N(\bm{k}_{2})\rangle-\langle N(\bm{k}_{1})\rangle\langle N(\bm{k}_{2})\rangle$
@@ IV.1 Definitions of the correlation functions
$+2\langle N(\bm{k}_{1})\rangle\langle N(\bm{k}_{2})\rangle\langle N(\bm{k}_{3})\rangle$ (S4)
and their integrals over zero-sum (quasi-)momentum modes
$\overline{G}_{c}^{(2)}(0)=$ $\sum_{\bm{k}\in\Omega_{k}}G_{c}^{(2)}(\bm{k},-\bm{k})$ (S5)
$\overline{G}_{c}^{(3)}(0)=$ $\sum_{\bm{k}_{1},\bm{k}_{2}\in\Omega_{k}}G_{c}^{(3)}(\bm{k}_{1},\bm{k}_{2},-\bm{k}_{1}-\bm{k}_{2})$ (S6)
@@
IV.2 Comparison between theory and experiment
@@
IV.3 Scaling of the correlation functions in the superfluid vs. in the Mott (or normal gas) limit
We discuss here the different scaling regimes of the momentum correlations predicted by the theory on quantum rotors. We restrict our attention to a system in the canonical ensemble at fixed particle number $N$, and at fixed density in the lattice $n_{\rm QR}=N/V$ where $V$ is the system volume, so that particle number and volume are proportional. In the superfluid regime we expect the two-mode correlations $G_{c}^{(2)}(\bm{k},-\bm{k})$ to be $\sim O(1)$, as predicted by Bogoliubov theory, and therefore we expect $\overline{G}_{c}^{(2)}(0)\sim O(N)$ since $V_{\Omega_{k}}\sim V$. This scaling is indeed observed in our quantum-rotor calculations. On the other hand we do not know of a simple theoretical prediction for $G_{c}^{(3)}(\bm{k}_{1},\bm{k}_{2},-\bm{k}_{1}-\bm{k}_{2})$ – indeed Bogoliubov theory predicts this type of correlation to be vanishing. Hence we resort to numerics in order to estimate the scaling of $\overline{G}_{c}^{(3)}(0)$, and we find from our quantum-rotor calculations that $\overline{G}_{c}^{(3)}(0)\sim O(N)$. This implies that three-mode correlations are much more tenuous than two-mode ones, namely $G_{c}^{(3)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\sim O(N^{-1})$ so that their integral over $O(N^{2})$ trios of modes is $O(N)$.
The scaling of the integrated correlations in the superfluid regime should be contrasted with that of a deep Mott-insulating or normal-gas regime, in which all quasi-momentum modes become statistically independent, modulo the conservation of the particle number. The population of each mode is Poisson-distributed, with average occupation $\langle N(\bm{k})\rangle=n$ where $n$ is the lattice filling; moreover one has $\langle N^{2}(\bm{k})\rangle=n(n+1)$ and $\langle N^{3}(\bm{k})\rangle=n^{3}+3n^{2}+n)$, according to Poisson statistics. A lengthy but straightforward calculation leads to the result
$G_{c}^{(2)}(\bm{k}_{1},\bm{k}_{2})$ $\to-\frac{n}{V-1}$
$G_{c}^{(3)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$ $\to\frac{2n}{(V-1)(V-2)}$ (S7)
(for inequivalent quasi-momenta $\bm{k}_{1}$, $\bm{k}_{2}$ and $\bm{k}_{3}$). This implies that
$\overline{G}_{c}^{(2)}(0)$ $\to-\frac{nV_{\Omega_{k}}}{V-1}$
$\overline{G}_{c}^{(3)}(0)$ $\to\frac{2n{\cal N}_{3,\Omega_{k}}}{(V-1)(V-2)}$ (S8)
@@
Fig. S3 shows results for the $\overline{G}_{c}^{(2)}(0)$ and $\overline{G}_{c}^{(3)}(0)$ correlations of the quantum-rotor model without any further renormalization factor. We observe that the asymptotic behavior predicted in the deep normal-gas limit at large interaction strength is correctly reproduced by our results. Interestingly, the integrated $\overline{G}_{c}^{(3)}(0)$ does not reach this limit monotonically from above, but it rather shows an intermediate minimum, corresponding roughly to the superfluid/normal-gas transition.
V Finite-temperature Bogoliubov theory
@@
VI Quantum-rotor model and Quantum Monte Carlo

### v2

@@ IV Scaling of the uncertainty on the value of $G^{(2)}_{c}({\bf 0})$ with interactions
This generic equation is valid for both Gaussian and non-Gaussian states. In the main text, we make use of this expression in the regime where Bogoliubov theory fails to show the presence of non-Gaussian correlations, i.e. $\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle_{c}\neq 0$.
@@
VI Three-body connected correlations in the experiment
As discussed in the main text, three-body connected correlations $\overline{G}_{c}^{(3)}(\mathbf{0})=\sum_{\bm{k}_{1},\bm{k}_{2}\in\Omega_{k}}G_{c}^{(3)}(\bm{k}_{1},\bm{k}_{2},-\bm{k}_{1}-\bm{k}_{2})$ are ideally suited to reveal momentum-correlated triplets induced by interaction. We have computed $\overline{G}_{c}^{(3)}(\mathbf{0})$ from the experimental data, using the definition
@@ VI Three-body connected correlations in the experiment
$+2\langle N(\bm{k}_{1})\rangle\langle N(\bm{k}_{2})\rangle\langle N(\bm{k}_{3})\rangle.$ (S9)
Unfortunately, the experimental signal-to-noise appears not sufficient to reveal the presence of a peak in $\overline{G}_{c}^{(3)}(\mathbf{0})/\Delta k^{3}$ over the range $\Delta k<0.2k_{d}$, which would be the analog of that observed in $\overline{G}_{c}^{(2)}(\mathbf{0})/\Delta k^{3}$ for two-body correlations (see Fig. S2). We illustrate this fact in figure S5 by plotting the measured $\overline{G}_{c}^{(3)}(\mathbf{0})$ at the value $U/J=10$.
We also compare $\overline{G}_{c}^{(3)}(\mathbf{0})$ with the measured values $\overline{G}_{c}^{(2)}(\mathbf{0})$. From our numerical calculations with quantum rotors, we expect their ratio to be maximum for $U/J=10$, with $\overline{G}_{c}^{(3)}(\mathbf{0})/\overline{G}_{c}^{(2)}(\mathbf{0})\sim 0.1$ (see fig. S7). Figure S5 illustrates that error bars on $\overline{G}_{c}^{(3)}(\mathbf{0})$ are too large to reveal such a signal.
Figure S5: Three-body connected correlation $\bar{G}^{(3)}_{c}(\bm{0})$ as a function of the size of the integration volume $\Delta k$ for $U/J=10$. The error bars correspond to 68%-confidence intervals obtained from a bootstrap method. Inset: zoom into the left panel with a comparison to $\bar{G}^{(2)}_{c}(\bm{0})$.
VII Correlation functions in momentum space and their scaling behavior
VII.1 Definitions of the correlation functions
In this work, we consider the following connected correlations of the (momentum) density operator $N(\bm{k})=a^{\dagger}_{\bm{k}}a_{\bm{k}}$: $G_{c}^{(2)}(\bm{k}_{1},\bm{k}_{2})=\langle N(\bm{k}_{1})N(\bm{k}_{2})\rangle-\langle N(\bm{k}_{1})\rangle\langle N(\bm{k}_{2})\rangle$ and $G_{c}^{(3)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$ as defined in Eq. S9, and their integrals over zero-sum (quasi-)momentum modes
$\overline{G}_{c}^{(2)}(\mathbf{0})=$ $\sum_{\bm{k}\in\Omega_{k}}G_{c}^{(2)}(\bm{k},-\bm{k})$ (S10)
$\overline{G}_{c}^{(3)}(\mathbf{0})=$ $\sum_{\bm{k}_{1},\bm{k}_{2}\in\Omega_{k}}G_{c}^{(3)}(\bm{k}_{1},\bm{k}_{2},-\bm{k}_{1}-\bm{k}_{2})$ (S11)
@@ VII.1 Definitions of the correlation functions
VII.2 Comparison between theory and experiment
All theoretical calculations are performed on cubic lattices with periodic boundary conditions. In the theoretical calculations, the wavevectors ${\bm{k}}$ are associated with discrete quasi-momentum modes living in the first Brillouin zone, and $N({\bm{k}})$ are the populations of this discrete set of modes.
@@ VII.2 Comparison between theory and experiment
VII.3 Scaling of the correlation functions in the superfluid vs. in the Mott (or normal gas) limit
We discuss here the different scaling regimes of the momentum correlations predicted by the theory on quantum rotors. We restrict our attention to a system in the canonical ensemble at fixed particle number $N$, and at fixed lattice filling $n_{\rm QR}=N/V$ where $V=L^{3}$ is the number of sites. In the superfluid regime we expect the two-mode correlations $G_{c}^{(2)}(\bm{k},-\bm{k})$ to be $\sim O(1)$, as predicted by Bogoliubov theory, and therefore we expect $\overline{G}_{c}^{(2)}(0)\sim O(N)$ since $V_{\Omega_{k}}\sim V$. This scaling is indeed observed in our quantum-rotor calculations. On the other hand we do not know of a simple theoretical prediction for $G_{c}^{(3)}(\bm{k}_{1},\bm{k}_{2},-\bm{k}_{1}-\bm{k}_{2})$ – indeed Bogoliubov theory predicts this type of correlation to be vanishing. Hence we resort to numerics in order to estimate the scaling of $\overline{G}_{c}^{(3)}(0)$, and we find from our quantum-rotor calculations that $\overline{G}_{c}^{(3)}(0)\sim O(N)$. This implies that three-mode correlations are much more tenuous than two-mode ones, namely $G_{c}^{(3)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\sim O(N^{-1})$ so that their integral over $O(N^{2})$ trios of modes is $O(N)$.
The scaling of the integrated correlations in the superfluid regime should be contrasted with that of a deep Mott-insulating or normal-gas regime, in which all quasi-momentum modes become statistically independent, modulo the conservation of the particle number. The population of each mode is thermally distributed so as to exhibit bunching, with average occupation $\langle N(\bm{k})\rangle=n_{\rm QR}$. Moreover one has $\langle N^{2}(\bm{k})\rangle=n_{\rm QR}(2n_{\rm QR}+1)$ and $\langle N^{3}(\bm{k})\rangle=6n_{\rm QR}^{3}+6n_{\rm QR}^{2}+n_{\rm QR}$, according to thermal statistics. A lengthy but straightforward calculation leads to the result
$G_{c}^{(2)}(\bm{k}_{1},\bm{k}_{2})$ $\to-\frac{n_{\rm QR}(n_{\rm QR}+1)}{V-1}$
$G_{c}^{(3)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$ $\to\frac{4n_{\rm QR}^{3}+6n_{\rm QR}^{2}+2n_{\rm QR}}{(V-1)(V-2)}$ (S12)
(for inequivalent quasi-momenta $\bm{k}_{1}$, $\bm{k}_{2}$ and $\bm{k}_{3}$). This implies that, for $n_{\rm QR}\gg 1$ (see Sec. IX)
$\overline{G}_{c}^{(2)}(0)$ $\to-\frac{n_{\rm QR}^{2}~{}V_{\Omega_{k}}}{V-1}$
$\overline{G}_{c}^{(3)}(0)$ $\to\frac{4n_{\rm QR}^{3}~{}{\cal N}_{3,\Omega_{k}}}{(V-1)(V-2)}$ (S13)
@@
Fig. S6 shows results for the $\overline{G}_{c}^{(2)}(0)$ and $\overline{G}_{c}^{(3)}(0)$ correlations of the quantum-rotor model without any further renormalization factor. We observe that the asymptotic behavior predicted in the deep normal-gas limit at large interaction strength is correctly reproduced by our quantum rotor results. As explained in the next section, this is to be expected Carcy et al. (2019), since in the deep normal-gas limit the momentum distribution of the quantum rotor model becomes a speckle pattern, whose exponential statistics Dainty (1975) is the same as the one of thermal Bose fields. Interestingly, the integrated $\overline{G}_{c}^{(3)}(0)$ does not reach this limit monotonically from above, but it rather shows an intermediate minimum, corresponding roughly to the superfluid/normal-gas transition.
VII.4 Ratio of three-body and two-body connected correlations
We discuss here the variation of the amplitudes of two-body and three-body connected correlations obtained with the quantum rotor results. In Fig. S7, we plot the ratio $G_{c}^{(3)}(0)/G_{c}^{(2)}(0)$ as a function of $U/2n_{\rm QR}J$. We identify two regimes, $U/2n_{\rm QR}J<4$ and $U/2n_{\rm QR}J\geq 4$.
Figure S7: Ratio $G_{c}^{(3)}(0)/G_{c}^{(2)}(0)$ versus $U/2n_{\rm QR}J$ obtained with the model of quantum rotors.
At small values of interactions $U/2n_{\rm QR}J<4$, both the amplitudes $G_{c}^{(3)}(0)$ and $G_{c}^{(2)}(0)$ increase with interaction (see main text). However, the increase of their ratio illustrates that $G_{c}^{(3)}(0)$ increases faster than $G_{c}^{(2)}(0)$. This finding is compatible with the microscopic picture depicted in Fig. 1 of the main text: as the condensate fraction decreases, the importance of the interaction term $V^{(3)}$ responsible for the emergence of correlated triplets grows.
At larger interactions $U/2n_{\rm QR}J\geq 4$, we find that the ratio $G_{c}^{(3)}(0)/G_{c}^{(2)}(0)$ is roughly constant. This indicates that the perturbative picture drawn in Fig. 1 of the main text does not apply to the equilibrium state at large interactions. This is compatible with the fact that, in this regime, the condensate fraction is small and that the various terms depicted in Fig. 1 have similar contributions.
VII.5 Behavior of the $G_{c}^{(2)}$ correlations at the superfluid/normal-gas transition
As seen in the previous subsection, $G_{c}^{(2)}(0)$ turns to negative values in the deep normal-gas limit, while it is positive in the deep superfluid regime because of Bogolyubov momentum pairing. How does $G_{c}^{(2)}(0)$ transition between these two regimes?
Fig. S8 shows quantum-rotor results for the critical behavior of $G_{c}^{(2)}(0)$ at the SF/NG transition driven by an increasing interaction at fixed temperature. The position of the transition is identified via the scaling of the condensate fraction. When examining the behavior of $G_{c}^{(2)}(0)$ in the vicinity of the transition one observes that it remains positive over the entire SF regime, and that it appears to vanish in the vicinity of the transition, to then turn negative in the NG phase. In particular the value of $U/(2Jn_{\rm QR})$ at which $G_{c}^{(2)}(0)$ crosses zero appears to move closer to the critical value as the system size increases, suggesting that the critical point may correspond to the point at which $G_{c}^{(2)}(0)$ changes sign.
Figure S8: Superfluid/normal-gas transition for the quantum-rotor model at $T/(2Jn_{\rm QR})=1$. Upper panel: scaling behavior of the condensate fraction at the SF/NG transition, identifying the critical point at $[U/(2Jn_{\rm QR})]_{c}$ Here $\beta=0.3485$ and $\nu=0.67155$ are critical exponents of the 3d XY universality class. Lower panel: behavior of $G_{c}^{(2)}(0)$ in the vicinity of the transition for $L^{3}$ cubic lattices. $G_{c}^{(2)}(0)$ is defined by integrating over a cubic corona with momentum modes $(2\pi/L)(n_{x},n_{y},n_{z})$ with $|n_{\alpha}|\geq L/2-2$ for a least one component $\alpha=x,y$ or $z$.
VIII Finite-temperature Bogoliubov theory
@@
IX Quantum-rotor model and Quantum Monte Carlo
@@
In particular, in the deep normal-gas limits the quantum-rotor variables $\phi_{i}$ become uncorrelated random angles, so that the quantum-rotor picture for the Bose field in momentum space
$b_{\bm{k}}\to\sqrt{\frac{n_{\rm QR}}{V}}\sum_{i}e^{i\bm{q}\cdot\bm{r}_{i}}e^{i\phi}=\frac{1}{\sqrt{V}}\sum_{i}\sqrt{n_{\rm QR}}~{}e^{i\theta_{i}}$ (S23)
becomes a sum of random complex numbers of fixed amplitude $\sqrt{n_{\rm QR}}$ and random phase $\theta_{i}=\bm{q}\cdot\bm{r}_{i}+\phi_{i}$. The momentum distribution is then the square modulus of this random complex variable, which, in the limit of a large system $V\gg 1$, has the same statistics of the intensity of a laser-speckle pattern Dainty (1975), namely
$P(N_{\bm{k}})\to P(|b_{\bm{k}}|^{2})=\frac{e^{-\frac{N_{\bm{k}}}{n_{\rm QR}}}}{n_{\rm QR}}~{},$ (S24)
namely the same statistics as that of thermal photons in the limit of a large average photon number, $n_{\rm QR}\gg 1$.